Given: C = 100 + 0.65Yd (where Yd = Y-T) I = 120-400i G = 200 T = 20 + 0.2Y Ms/P = 200 Md/P = 50+0.5Y-600i Where: C = Consumption Y = Income I = Investment G = Government spending T = Taxes i = interest rate Ms/P = RealMoney Supply Md/P = Real Demand for Money (a) Derive the IS and LM curves (10 Marks) (b) Obtain the equilibrium level of: i. Income (4 marks) ii. and consumption


1
Expert's answer
2021-05-10T16:41:50-0400

a.

The equation of IS curve is derived as follows:

"Y= AE"

"Y=C + I + G"

"Y= 100 + 0.65 Yd + 120 \u2212 400i + 200"

"Y=100 + 0.65 (Y \u2212 T) + 120 \u2212 400i + 200"

 "Y=100 + 0.65 (Y \u2212 20 \u2212 0.2Y) + 120 \u2212 400i + 200"

"Y= 847.91 \u2212 833.33 i ... (1)"


 The equation of LM curve is derived as follows:

"\frac{Md}{P} =\frac{ms}{p}"

"50 + 0.5Y \u2212 600i = 200"

"Y = 300 + 1200i ... (2)"


b.

The equilibrium level of income and interest rate is calculated using the equation (1) and (2) as follows:

"300+1200i=847.91-833.33i"

"1200i+833.33i=847.91-300"

"2033.33=547.91i"

"i= 0.2694"


"Y=300+1200i"

"Y=300+(1200\times0.2694)"

"Y =623.36"


The equilibrium level of consumption is calculated as follows:

"C= 100 + 0.65 (Y \u2212 T)"

"=100 + 0.65 (Y \u2212 20 \u2212 0.2Y)"

"=87 + 0.52Y"

"= 87 + 0.52 (623.36)"

"= 411.14"